Gundam 0083 Stardust Memory - Episode 13: The Storm That Runs Through Page 1

Gundam 0083 Stardust Memory Episode 13: The Storm That Runs Through (English Title: Men of Destiny)

Translated By Tonghyun “Vajra” Kim (tkim@sdcc3.ucsd.edu) Edited by Noel Gamboa (egamboa@sdcc13.ucsd.edu) OPERATION X Version 1.0 11/22/92

Gato, screaming in anger, punches a hole in the Gwaden’s bridge with his Neue Ziel!

Title

The Storm That Runs Through

Sinapus gets new orders from Basque Ohm.

Sinapus

What the hell!? Did you say protect the Solar System with them?

Basque Ohm

That’s right. As of now, the Cima Fleet is cooperating with us.

Sinapus

This is just like a military dictatorship!

Basque Ohm

You dare defy the order?

Uraki hears this same news…

Scott

Lt. Uraki, you can’t catch up with Gato from that position. Please retreat immediately!

Kou

What did Lt. Burning die for? And Kelley as well!

Uraki kicks in the thrusters of the slightly damaged Dendrobium.

Scott

Lt. Uraki, where are you going?

Meanwhile, Gato blasts towards the Solar System, the only thing standing in the way against his mission’s success.

Gato

210 more minutes. I won’t let those mirrors…

Bridge operator

Enemy mobile armor approaching. Distance 12000!

Basque Ohm

Damn that Cima, she blew it! Buy us time until the firing of the Solar System!

The Federation ships open fire at the approaching MA. Meanwhile, leaves the Gwaden on her new MS, the Gerbera Tetra. Quite naturally, she’s pissed at Gato for ruining her plans and almost killing her by decompressing the bridge.

Cima

You’ll pay for ruining my gamble!

Cima fires at the Gwaden’s rear hangar, triggering a series of explosions which envelop the Gwaden. While all hell breaks out in front of the mirror array, the control ship staff make final preparations for firing.

Operator 1

81 percent completed. We’re behind schedule!

Operator 2

F1, F5 and F6. Focusing standby.

Operator 3

L15, hurry the array check!

Gato rockets towards the mirrors.

Gato

Damn… if I can just destroy the control ship…

Cima approaches her flagship, the Lili Marleen.

Cima

Don’t light the guide beacon! Do you want to get killed or what?

Cima

The sensors! Above!

A bolt of light from 12 o’clock high strikes the Lili Marleen, destroying it.

Cima

I knew it.

A crazed Kou heads for the Cima Fleet vessels protecting the Solar System.

Scott

Lt. Uraki, the Solar System will fire at any moment! Lt. Uraki!

Gato, meanwhile, is cutting down some Fed ships.

Gato

Get out of my way!

Basque Ohm’s ship. He’s worried about the rampaging Neue Ziel.

Bridge operator

Distance to space colony: 4000. 70 seconds till firing of the Solar System.

Basque Ohm

We can see the colony with our eyes. We’ve waited long enough! Fire!

Basque Ohm’s staff

But we have to wait 30 seconds more…

Basque Ohm

I don’t care! Just do it!

Gato makes contact with the mirror surface, and begins to search for the control ship.

Gato

Huh? That’s the one?

The mirrors align and fire a bolt of pure light towards the falling colony. Gato levels the Neue Ziel’s beam guns on the control ship and fires!

Gato

God help us!

The colony is enveloped in light and smoke. The control ship, however, explodes.

Bridge operator

Colonel! The control ship!

Basque Ohm

What!?

The colony appears from behind billows of smoke. It is now directly in front of the mirrors and the defending ships, both Federation’s and Cima Fleet’s.

Basque Ohm

It survived…

Basque Ohm

What are you doing? Evasive! Evasive now!

Ship captain

Turn away. Turn away!

Ships collide in space, and the colony pushes through the array, scattering the pieces.

Gato

No one can stop it now.

Gato smiles briefly.

Gato

All that’s left…

Albion

Simmone

The colony has entered its final deceleration course. 151 minutes before collision.

Sinapus

What’s Uraki doing? He still hasn’t returned?

Mora

Captain! Nina went out from the closed port deck!

Sinapus

What?

Nina flies the Core Fighter into the colony.

Keith

Nina? Is she headed for the colony?

Kou drops in on his friends from Cima Fleet.

Uraki

You bastards!

Gelgoog pilot

Madame Cima! Please retreat!

Cima

Where the hell can we retreat to?

Cima

Explosion Inducement Chain! He stopped moving?!

Kou destroys a ship using the Chain.

Cima

You bastard, whose side are you on?

Cima rains bullets on the Dendrobium, but Kou slams his cannon into the Gerbera Tetra.

Cima

What?

Cima’s capsule experiences decompression. Cima is dying, but Kou decides to blast her to molecules anyway. When Cima is space dust, Kou remembers his archrival…

Kou

Gato!

Gato has docked the Neue Ziel outside the colony. Now, inside the colony’s control room, he sets the program to adjust the colony’s trajectory.

Gato

With the success of this operation, there will be those who will follow in our path.

Gato

Nina Purpleton.

Nina

Please, Gato. Isn’t this enough? Please stop it!

Gato

It’s not over yet! I will complete Stardust.

Nina

This colony will definitely fall on Earth. It doesn’t have to fall on Jaburo.

Gato

We’re not dropping it on Jaburo. You can’t understand.

Nina

Why? Why did you appear before me again? After suddenly disappearing that day…

Gato

I wanted you to forget everything (about those days)… when I stayed on the moon, anxiously waiting for the time to come.

Nina

I tried to forget everything! No, I had forgotten them!

Gato

I have devoted my body and soul to the resurrection of Jion. You might be the true witness of Stardust.

Nina finds Gato’s gun and threatens to shoot. A shot is fired, however, by someone else – Kou, who has entered the control room as well. Gato still manages to pull the lever, activating the program.

Nina

Gato! Hang on!

Kou

Nina!

Nina

Please stop, Kou. There’s no longer any reason for you to fight.

Kou

Get away from him! Don’t you know what he’s done?

Nina

I had to witness the battle between the two of you, until one of you would die… that’s why I was praying that things don’t end like this!

Kou

Nina, were you lying about your feelings for me? Tell me!

Nina

That’s not true…

Nina

Gato!

Gato pushes Nina aside and stands erect before Kou. Kou’s gun remains pointed at Gato, but Nina suddenly points hers at Kou.

Nina

Kou, please stop.

Kou

Nina… This can’t be real. Nina! He dropped the colony. He stole Gundam Unit 2!

Nina

Kou, it’s not about those things…

Kou

Nina, come back! Please come to me!

Nina and Gato disappear behind a door.

[AXIS ADVANCE FLEET]

Operator

Rescuing soldiers is the first priority! Step up the retrieval operation! Abandon units that have been damaged severely.

Axis bridge officer

The colony has completed the final course adjustment. It’s perfect, sir.

Hustler

Delaz, the Stardust has been born. You made it…

Jion crew

Commander. The Federation pursuit fleet from the moon has arrived.

Konpei base commander (pursuit fleet commander)

Inform them (Basque Ohm’s Fleet) of the retreat time in the agreement again, to attack the Axis Fleet as soon as the duration of their stay expires.

Nina patches Gato’s uniform.

Gato

I have a rough idea of what happened between you and him. If that’s the case, I want you to leave me even more.

Nina

At any rate, I had no choice but to do this.

Gato

Nina… I’m sorry.

Gato punches Nina, who loses consciousness. The Neue Ziel then leaves the colony clutching the Core Fighter’s escape hatch.

Karius

Commander, we’ll be attacked from both sides by the enemy from the moon. Please escape with us. We don’t have time!

Gato

Karius…

Kou’s Dendrobium Stamen exits the colony.

Scott

Lt. Uraki. The colony has begun to disintegrate. Please escape quickly. Can you hear me? We did our best.

Uraki

This is Uraki. I just left the colony.

Scott

Lieutenant, you’re safe! Please join with the Albion immediately.

Kou

W - What?

Outside, the Neue Ziel awaits Kou Uraki.

Scott

The Albion’s at point 771, 8000 behind the colony. We can’t keep up any more. Please hurry!

Gato

If you weren’t a part of the rotten Federation, there’s no need for you to go through all this anguish…

Kou activates his weapons system.

Scott

Lieutenant, what’s wrong? Lieutenant!

Kou

He’s been waiting… for me?

Several Rick Dom mobile suits, led by Karius, head for the Axis forces. Karius’s suit is clutching the pod where Nina is.

Karius

Why does she have to go through all this? That light!

Basque Ohm has partially rebuilt the mirror array with less than 50 percent of the original surface area.

Operator

The enemy mobile armor is currently circling the Earth orbit and will soon appear from the other side of Earth.

Operator

Due to the loss of the control ship, the efficiency of the second blast will be under 25 percent.

Basque Ohm’s staff

Colonel. The enemy seems to be engaged in a battle with Gundam Unit 3…

Basque Ohm

Shut up! I won’t be able to contain myself until I blast ’em!

The battle between the Delaz Fleet and the Federation ships continue.

Scott

Engine output has dropped! And the second propellant tank won’t open. Direct hit to starboard. Main cannon is down!

Simmone

Battle speed reduced to 25.

Sinapus

We paid a heavy price… We’ve gone as far as we can go. Retreat from the battle zone!

Kou and Gato dogfight. Gato drops the NZ’s fuel tanks, turns, and severs the Dendrobium’s right arm with a beam blast. Kou activates the left arm’s beam saber, but Gato wrestles it with his NZ’s right arm [the only one left].

Gato

You don’t have any weapons left!

The arm’s control line is cut.

Kou

Behind?

Gato

Too slow!

Four small, grappling arms pop out of the NZ! Gato drops over the Dendrobium and the grappling arms quickly embrace the Dendrobium’s torso, preventing MS/armed base separation.

Gato

What next? You can’t even separate.

The Solar System begins to fire…

Kou

What the…

Basque Ohm’s staff

Please wait! Our advance fleet is right in front of the target! Colonel!

Kou and Gato are enveloped by white light.

On Jaburo, site of the Federation’s headquarters, two soldiers are escorting General Kowen, possibly to a military prison or a similar facility.

Federation officer

Hey, the evacuation’s been cancelled.

Federation soldier

The colony won’t fall here!

Federation soldier

What’s that?

The glowing colony passes over Jaburo.

Kowen

This incident will change history…

space

Kou

Gato… He left without finishing me off…

Axis Advance Fleet; main ship.

Operator

Walfisch! Do you read me? Walfisch!

Operator

Commander Gato, we don’t have time.

Hustler approaches the weeping Nina.

Hustler

You’re Nina, aren’t you? Look.

Nina

The colony… is falling on Earth.

Hustler

That’s the shining light from the souls of men.

Gato, in a partially burned Neue Ziel, leads the Delaz Fleet survivors to where Axis forces can pick them up. A large Federation contingent, however, is in the way.

Gato

Come on, let’s go.

Basque Ohm’s staff

Remaining units of the Delaz Fleet, listen. You have nowhere to return now. Please surrender at once. We know that you no longer have any combat potential. There is no need to throw away your lives.

Gato

Understand? Break through and join the Axis Fleet, even if only one of us makes it. So that you can tell the truth about our fight to future generations!

Basque Ohm’s staff (to axis forces)

The duration of your fleet’s stay has expired. Leave at once! Or do you wish to engage in battle with us?

Hustler

We’ll conclude our retrieval operation.

Nina

Five more minutes. Please wait five more minutes!

Hustler

Change heading. Set course for Axis.

Hustler

I want you to make a free choice.

Nina

I beg you!

Gato leads the charge against the Federation blockade. Despite his injury and the damage to the Neue Ziel, he destroys several ships. In one final act of defiance, he slams his mobile armor into a ship, destroying it and the Neue Ziel, along with himself. From the main Axis ship, a small vessel leaves for Earth.

Kou’s trial [for disobeying orders/stealing Unit 3]

Federation prosecutor

Keeping quiet won’t do you any good! Answer me! Are you deaf?

Federation judge

Lt. Junior Grade Kou Uraki, one year in prison.

On the same day as Lt Uraki’s trial, Captain E Sinapus was sentenced to death, according to the LD liner notes.

Basque Ohm

Look back. Those optimists who dreamed about the establishment of the Earth Sphere… have caused this incident! The Delaz Fleet is just an example. You don’t need to look at the colony fall accident, which inflicted severe damage to the granary of North America three days ago, to realize that our Earth is constantly exposed to various dangers.

Images flash… Anaheim’s vice-CEO commits supposedly suicide… the Albion bridge crew join the Titans…

Basque Ohm

Earth, symbol of the Universe, we are born to protect her! Earth! In order to return her true power… to those who truly deserve it, the Titans will stand!

Text

The Gundam Development Plan under the supervision of General Kowen was removed from all records.

Text

Gundam GP01-fb GP02A GP03S Deleted from all records.

Text

Consequently, Kou Uraki’s criminal record was erased as well.

Three months after his trial, Kou is reassigned to the Federation Oakly [Oakley?] Base in Kansas, North America.

Kou

Gato…

Kou sees a Gelgoog used by the Federation for testing. The Gelgoog flashes a ”peace” sign.

Kou

Keith!

Kou spots a jeep with Mora and Nina.

Kou

Nina… Opening Song:

Evergreen (long version)

Alone in the deep darkness, You lose your way to the night Continuing to search for the memory of a shining dawn. Sometimes, destiny is a wind that tests the heart Love is the sky that covers everything. Showing the vision of tomorrow.

DREAMER. Burning heartbeats That’s what we have to bring back.

DREAMER. The frozen dream Is beginning to shine again.

EVERGREEN.

The arms that are empty now are endlessly looking for someone you can share your life with.

DREAMER What we need to do is to overcome The burden of the time.

DREAMER And breath new life To the frozen dream.

EVERGREEN

DREAMER. Burning heartbeats That’s what we have to bring back.

DREAMER What we need to do is to overcome The burden of the time.

Planning:  Sunrise
Original Concept:  Yadachi Hajime

Tomino Yoshiyuki

Executive Producer:  Ueda Masuo (Sunrise)

Takanashi Minoru (Bandai)

Executive Animation Director / Character Designer:

Kawamoto Toshihiro

Executive Mechanical Animation Director:

Sano Hirotoshi

Mechanical Styling:  Kawamori Shoji
Mechanical Designer:  Katoki Hajime

Akitaka Mika (RED) Ishizu Yasushi

Director of Art:  Higashi Junich
Director of Photography: Okui Atsushi
Music by  Hagita Mitsuo
Sound Director:  Uragami Yasuo

OP Men of Destiny sung by MIO composed by Matsuhara Miki

[First OP The Winner sung by Matsuhara Miki]

ED Evergreeen sung by MIO composed by Suzuki Kisaburo

[First OP Magic sung by Jacob Wheeler]

Director:  Imanishi Takashi

Kase Mitsuko

Screenplay:  Ookuma Asahide

Gobu Fuyushi Endo Akinori Takahashi Ryosuke

Cast

Anavel Gato  Ootsuka Akio
Kou Uraki  Horikawa Ryo
Nina Purpleton  Sakuma Rei
Eyphar Sinapus  Ootsuka Chikao
Chack Keith  Yamada Yoshiharu
Mora Bashit  Ikura Kazu
South Burning  Sugahara Masashi
Aiguille Delaz  Kobayashi Kiyoshi
Cima Garahau  Mashiba Mari
Karius  Tobida Nobuo
Wille Gradol  Inaba Minoru
Kelley Layzner  Genda Tessho
Basque Ohm  Gouri Daisuke

Jamitov Hyman  Nishimura Tomomichi
Green Wyatt  Tanaka Hideyuki
John Kowen  Watabe Takeshi
Bernard Monsha  Cha Furin
Alpha Bait  Toya Kouji
Chap Adel  Mikimoto Masayuki
Peter Scott  Kikuchi Masami
Simmone Jacqueline Araki Kagae
Ivan Basrof  Kishino Kazuhiko
Paula Guillish  Matsui Naoko
Lucette Audevie  Katsuo Masako
Latuera Chapra  Kobayashi Yuuko

Yuli Hustler

Hamaan Khan  Sakakibara Yoshiko

Animation Director (Eps. 13)

Kawamoto Toshihiro

Mechanical Animation Director (Eps. 13)

Sano Hirotoshi

Key Animators:  Katoki Hajime  Sano Toshihiro
Kawamoto Toshihiro  Osaka Hiroshi
Yoshida Tooru  Yamashita Takehito
Horikawa Kouichi  Naka Narifumi
Ito Kouji  Minahiro Kazumi
Seo Yasuhiro  Sugano Hiroki
Tsunoda Katsutoshi  Sugiura Kouji

Animation Check:

Animator:  Watanabe Keiichi  Yamashita Hiroshi
Samura Yoshikazu  Akasaka Rieko

Yamaguchi Yoshitaka

Color supervision/establishment:

Nakayama Shihoko

Coloring:  KYUMA  Yamada Shin
Matsumoto Masae  Studio Bogey
Kanezawa Keiko  Koyama Miyuki

Suzuki Animation Planning

FAI  Studio Ado

Production Planning:

Production Assistant:

Assistant Producer:  Minami Masahiko (Sunrise)

Terada Takehito

Production Manager

Storyboard:  Imanishi Takashi

Kase Mitsuko Watanabe Shinichiro

Producer:  Ookuma Asahide

Imanishi Takash Kase Mitsuk Watanabe Shinichiro

Production:  Bandai/Sunrise
Copyright:  Sunrise

THE END